# JEL classification: C31, J24

### v1

@@
[…]
Peer influence on effort devoted to some activity is often studied using proxy variables when actual effort is unobserved. For instance, in education, academic effort is often proxied by GPA. We propose an alternative approach that circumvents this approximation. Our framework distinguishes unobserved shocks to GPA that do not affect effort from preference shocks that do affect effort levels. We show that peer effects estimates obtained using our approach can differ significantly from classical estimates (where effort is approximated) if the network includes isolated students. Applying our approach to data on high school students in the United States, we find that peer effect estimates relying on GPA as a proxy for effort are 40% lower than those obtained using our approach.
@@ 1   Introduction
In recent years, there has been a growing interest in the impact of peers on educational outcomes (Sacerdote, 2011; Epple and Romano, 2011). As peer effects on students’ academic effort may involve a social multiplier effect, understanding whether students are influenced by their friends and the size of this influence is crucial for evaluating policies aimed at improving academic achievement (Manski, 1993). However, estimating peer effects based on academic effort presents challenges, since effort is generally unobserved. Consequently, many empirical studies resort to using grade point average (GPA) as a proxy variable for effort. While microfounded models exploring the impact of peer interactions on academic achievement are often focused on effort (e.g., see Calvó-Armengol et al., 2009; Arcidiacono et al., 2012; Fruehwirth, 2013, 2014; Hong and Lee, 2017), GPA is typically employed in the empirical analysis because measures of student effort are unavailable. Yet, this approximation overlooks that GPA is not solely influenced by effort but also by many other factors, including unobserved student and school characteristics.
@@
In this paper, we investigate the implications of using GPA as a proxy for academic effort on the estimation and interpretation of peer effects. We develop a structural model of educational effort with social interactions, in which students decide on their academic effort while taking into account that of their peers (friends), assuming complementarity between students’ and peers’ efforts. We also explicitly model the production function of academic achievement (GPA), which includes effort as a key input. Without using a proxy for student academic effort, we show that the model allows for identifying peer effects on academic effort. Furthermore, we demonstrate that the size of the effect may differ from that estimated using GPA as a proxy for effort, because the proxy-based approach suffers from an issue akin to an omitted variables problem. We derive a reduced-form equation for GPA, which differs from the classical linear-in-means peer effects specification, that highlights the need to control for two types of unobserved GPA shocks to disentangle peer effects on academic effort from other common effects captured by GPA. First, one needs to account for common shocks that directly influence GPA, such as improvements in teaching quality, irrespective of the effort level. These shocks result in a GPA increase for the same level of effort and do not involve a social multiplier effect. Second, one needs to account for common shocks affecting students’ preferences, such as increasing motivation to value academic achievement through information, which influence both academic effort and GPA, and may have social multiplier effects among students with friends.
@@
We demonstrate that approximating student effort with GPA may result in biased estimates of peer effects when some students in the network do not have friends. This is because students who have friends are affected differently by the two types of GPA shocks mentioned above compared to those without friends. Standard approaches using GPA as a proxy fail to differentiate between these two types of shocks, leading to biased estimates of peer effects when the network includes students without friends. However, we show that in networks without isolated students, the standard model does not produce biased estimates of peer effects. The key difference between the reduced-form of the standard model and our framework is that while the standard model has a single intercept for each school, our approach accounts for unobserved school-level heterogeneity based on whether a student has friends or not. This amounts to introducing two types of fixed effects per school, one for students with friends and one for isolated students.
We illustrate the importance of this distinction through both a Monte Carlo simulation study and an empirical application using the National Longitudinal Study of Adolescent to Adult Health (Add Health) data. [^1] The Add Health dataset comprises 22% of students without friends, including 11% who are not fully isolated, in the sense that they are nominated as friends of others. Other studies reporting social network data from educational settings with isolated students include those by Alan et al. (2021), Conti et al. (2013), and Boucher et al. (2021). Estimating peer effects through our proposed method suggests that increasing the average GPA of peers by one point leads to a 0.856 point increase in students’ GPA. In contrast, the standard linear-in-means model using GPA as a proxy for student academic effort estimates this effect at 0.507. This substantial difference highlights that failing to account for two types of shocks at the school level may yield substantially biased estimates. The intuition behind this bias is that the standard approach misidentifies direct shocks to GPA that do not influence effort as effort shocks, thereby overestimating their impact through a social multiplier effect. Importantly, we also find that other plausible approaches, such as estimating a model that accounts for only one type of school-level shock and incorporating a dummy variable for isolated students, fail to rectify the bias, as does excluding isolated students from the sample.
@@
Our econometric model incorporates two types of school-level unobserved shocks to GPA, which renders the well-known reflection problem more complex. [^2] The reflection problem arises when one cannot disentangle endogenous peer effects from exogenous contextual peer effects (Manski, 1993) In the case of the standard linear-in-means model, Bramoullé et al. (2009) provide straightforward conditions relating to the network structure under which the reflection problem is resolved. We extend their identification analysis to our framework; our main condition for identification requires that the network must include at least two students separated by a path of distance three, which is slightly stronger than the condition in Bramoullé et al. (2009).
@@
We also extend our analysis to the case of endogenous networks. Network endogeneity can occur because we do not observe certain student characteristics, such as intellectual quotient (IQ), that may influence both students’ likelihood to form links with others and their GPA. We control for these unobserved characteristics using a two-stage estimation approach. Our method is nonparametric as in Johnsson and Moon (2021). We do not impose a specific parametric restriction in the relationship between the unobserved characteristics and GPA. Our approach is similar to generalized additive models (GAM), which are widely employed in the nonparametric regression literature (see Hastie, 2017).
@@
Our structural model and econometric approach can also be applied to study peer influence on other outcomes that depend on exerted effort. An example is the body mass index (BMI), which cannot be directly chosen (e.g., Fortin and Yazbeck, 2015). People need to exert effort, such as developing healthy diet habits or engaging in physical exercise to improve their BMI. Peer influences are more related to effort than BMI. Another example is peer effects on a worker’s effort (e.g., Mas and Moretti, 2009; Cornelissen et al., 2017). The observed outcome is generally worker’s productivity, whereas peer effects take source in effort.
@@
There is a large literature studying social interactions both theoretically and empirically (Durlauf and Ioannides, 2010; Blume et al., 2011). We follow the games in networks approach to the analysis of social interactions (see Jackson and Zenou, 2015, for a comprehensive overview of this literature). Ballester et al. (2006) analyze a noncooperative game with linear-quadratic utilities and strategic complementarities, in which each player decides how much effort they exert. Applying a similar framework to an education application, Calvó-Armengol et al. (2009) use GPA as a proxy for the exerted effort. We contribute to this strand of the literature by explicitly modelling the production function that captures how effort, along with other factors, translates into GPA, while distinguishing between unobserved shocks that directly impact GPA without affecting effort, from shocks that impact both academic effort and GPA. This leads to a reduced-form equation for GPA that differs from the standard linear-in-means peer effects specification.
@@
This paper makes a methodological contribution to the extensive empirical literature on peer effects on educational outcomes (Sacerdote, 2011; Epple and Romano, 2011). We show both analytically and through an empirical application using AddHealth data that approximating student effort with GPA may result in biased estimates of peer effects, when some students in the network lack friends. Since isolated students are a common feature in many social network datasets, this finding highlights the limitations of using proxy variables, such as GPA, for estimating peer effects reliably.
@@
Our paper also contributes to the econometric literature on peer effects (De Paula, 2017; Kline and Tamer, 2020) in two ways. First, a key challenge in this field is the reflection problem (Manski, 1993). A recent wave of papers have addressed this issue by imposing conditions on the network structure (Bramoullé et al., 2009; De Giorgi et al., 2010). [^3] Other studies address the reflection problem using group size variation (Davezies et al., 2009; Lee, 2007) or imposing restrictions on the error terms (Graham, 2008; Rose, 2017). For an overview of this literature, see Bramoullé et al. (2020). Our contribution is to study the reflection problem in a setting where the GPA is influenced by various types of common shocks at the school level, and where students may have no peers. As in Bramoullé et al. (2009), our main identification condition involves the network structure and can be readily tested in empirical applications. Our second contribution lies in addressing the network endogeneity issue. We allow for unobserved attributes to influence students’ likelihood of forming links with others and their GPA (Goldsmith-Pinkham and Imbens, 2013; Hsieh and Lee, 2016; Johnsson and Moon, 2021). Unlike many studies that impose a strong parametric restriction between unobserved attributes and the peer effect model, we adopt a nonparametric approach to connect GPA to these attributes. Our approach is similar to the control function method used by Johnsson and Moon (2021), but it is adaptable to complex models. Specifically, we employ cubic B-splines, which are commonly used in Generalized Additive Models (GAM), to establish a smooth function linking the unobserved attributes and GPA (Hastie, 2017).
@@
The remainder of the paper is organized as follows. Section 2 presents the microeconomic foundations of the model using a network game in which students decide their academic effort. Section 3 describes the econometric model and addresses the identification and estimation of the parameters. Section 4 presents our empirical analysis using Add Health data. Section 5 provides an extension of our framework to endogenous networks. Section 6 concludes this paper.
@@ 2   The Structural Model
We consider $S$ independent schools and denote by $n_{s}\geq 3$ the number of students in the $s$-th school, $s\in\{1,\dots,S\}$. Let $n$ be the total number of students, that is, $n=\sum_{s=1}^{S}n_{s}$. Each student $i$ in school $s$ has a GPA denoted by $y_{s,i}$ and observable characteristics represented by a $K$-vector $\mathbf{x}_{s,i}$. Students interact in their school through a directed network that can be represented by an $n_{s}\times n_{s}$ adjacency matrix $\mathbf{A}_{s}=[a_{s,ij}]_{ij}$, where $a_{s,ij}$ = 1 if student $j$ is $i^{\prime}s$ friend and $a_{s,ij}=0$ otherwise. We assume that $a_{s,ii}$ = 0 for all $i$ and $s$ so that students cannot interact with themselves. In addition, we only consider within-school interactions; students do not interact with peers in other schools (see Calvó-Armengol et al., 2009). We define the social interaction matrix $\mathbf{G}_{s}=[g_{s,ij}]_{ij}$ as the row-normalized adjacent matrix $\mathbf{A}_{s}$, that is, $g_{s,ij}=1/n_{s,i}$ if $j$ is a $i$’s friend and $g_{s,ij}=0$ otherwise, where $n_{s,i}$ is the number of friends of student $i$ within school $s$.
Student $i$ chooses their effort $e_{s,i}$, which in turn affects their GPA. More precisely, GPA is a function of student effort $e_{s,i}$, observable characteristics $\mathbf{x}_{s,i}$, and a random term $\eta_{s,i}$ that captures unobservable characteristics. [^4] Whether or not the random term $\eta_{s,i}$ is observed by student $i$ or their peers is inconsequential for the analysis of the game. Following Fruehwirth (2013) and Boucher and Fortin (2016), we posit that this relationship is given by the production function:
$y_{s,i}=\alpha_{s}+\delta e_{s,i}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\theta}+\eta_{s,i},$ (1)
where $\delta>0$ and $\alpha_{s}$ and $\boldsymbol{\theta}$ are unknown parameters. The parameter $\alpha_{s}$ captures unobserved school-level GPA shifters, such as teacher quality, operating as fixed effects. The linear production function is somewhat restrictive and can be generalized to a nonlinear or nonparametric production function (e.g., see Fruehwirth, 2013). However, identification of the resulting econometric model can be intractable, especially when the model allows for unobserved school heterogeneity.
The effort exerted and the GPA obtained provide students with a benefit that is captured by a payoff function, which, as in Calvó-Armengol et al. (2009) and Boucher and Fortin (2016), takes a linear-quadratic form:
@@
where $\mathbf{X}_{s}=(\mathbf{x}_{s,1},\dots,\mathbf{x}_{s,n_{s}})^{\prime}$, $\mathbf{g}_{s,i}$ is the $i$-th row of $\mathbf{G}_{s}$, $\mathbf{e}_{s,-i}=(e_{s,1},\dots,e_{s,i-1},e_{s,i},\dots,e_{s,n})^{\prime}$, $\mathbf{e}_{s}=(e_{s,1},\dots,e_{s,n_{s}})^{\prime}$, the term $\mathbf{g}_{s,i}\mathbf{e}_{s}$ is the average effort of peers, and $c_{s}$, $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$ are unknown parameters. The parameter $\lambda$ captures endogenous peer effects. The payoff function (2) encompasses two components: a private sub-payoff and a social sub-payoff. The term $(c_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\varepsilon_{s,i})$ represents the benefit enjoyed per unit of GPA achieved, where $\varepsilon_{s,i}$ is the student type (observable by all students). This benefit accounts for student observed heterogeneity, as it depends on $\mathbf{x}_{s,i}$ and peer group average characteristics $\mathbf{g}_{s,i}\mathbf{X}_{s}$ termed contextual variables (see Manski, 1993). The benefit also accounts for school unobserved heterogeneity through the parameter $c_{s}$. The second term of the private sub-payoff, $e_{s,i}^{2}/2$ reflects the cost of exerting effort. The social sub-payoff $\lambda e_{s,i}\mathbf{g}_{s,i}\mathbf{e}_{s}$ implies that an increase in the average peer group’s effort $\mathbf{g}_{s,i}\mathbf{e}_{s}$ influences student $i$’s marginal payoff if $\lambda\neq 0$. When $\lambda>0$, the payoff function (2) implies complementarity between students’ and peers’ efforts, whereas $\lambda<0$ indicates substitutability in efforts. [^5] An alternative specification considers the social-payoff as $-\frac{\lambda}{2}(e_{s,i}-\mathbf{g}_{s,i}\mathbf{e}_{s})^{2}$, which represents a social cost depending on the gap between the student’s effort level and the average peer effort. This specification leads to conformist preferences when $\lambda>0$. Our approach can also be extended to this alternative specification.
@@
The parameters $\alpha_{s}$ and $c_{s}$ capture different unobserved shocks at the school level, and are conceptually different in terms of their policy implications. $\alpha_{s}$ captures unobserved shocks on GPA that do not affect student effort. These shocks, such as variation in teaching quality and school management, directly impact GPA irrespective of student effort. [^6] This reasoning holds because GPA is unbounded. An increase in $\alpha_{s}$ necessarily results in a higher GPA, and thus a higher payoff, regardless of the effort level. If we were to consider a framework where GPA is bounded, an increase in $\alpha_{s}$ may decrease the effort for students nearing the upper bound of GPA (e.g., see Fruehwirth, 2013). However, an increase in $\alpha_{s}$ does not have the same implication as an increase in $c_{s}$, which is the key factor in our framework. On the other hand, $c_{s}$ captures shocks on student preferences, particularly on the marginal payoff. For instance, interventions aimed at making students more aware of the returns to academic achievement could influence the marginal payoff. Such a shock can be captured by $c_{s}$ and would influence effort and consequently GPA through equation (1). We show that $\alpha_{s}$ and $c_{s}$ do not impact GPA in the same way (see Section 3.1).
@@
By substituting $y_{s,i}$ from Equation (1) into Equation (2), we obtain a payoff function that does not depend on GPA (see Appendix A.1). This new payoff function defines a static game with complete information, in which students simultaneously choose their effort levels to maximize their payoff. The best response function for the students is given by:
$e_{s,i}=\delta c_{s}+\lambda\mathbf{g}_{s,i}\mathbf{e}_{s}+\delta\mathbf{x}_{s,i}^{\prime}\boldsymbol{\beta}+\delta\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\delta\varepsilon_{s,i}.$ (3)
@@
The best response function (3) in matrix form can be expressed as $\mathbf{e}_{s}=\delta c_{s}\mathbf{1}_{n_{s}}+\lambda\mathbf{G}_{s}\mathbf{e}_{s}+\delta\mathbf{X}_{s}\boldsymbol{\beta}+\delta\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\delta\boldsymbol{\varepsilon}_{s}$, where $\boldsymbol{\varepsilon}_{s}=(\varepsilon_{s,1},\dots,\varepsilon_{s,n_{s}})^{\prime}$ and $\mathbf{1}_{n_{s}}$ is an $n_{s}$-vector of ones. A solution of this equation in $\mathbf{e}_{s}$ is a Nash equilibrium (NE) of the game. As $\mathbf{G}_{s}$ is row-normalized, the NE is unique under Assumption 2.1 and can be expressed as $\mathbf{e}_{s}=(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})^{-1}(\delta c_{s}\mathbf{1}_{n_{s}}+\delta\mathbf{X}_{s}\boldsymbol{\beta}+\delta\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\delta\boldsymbol{\varepsilon}_{s})$, where $\mathbf{I}_{n_{s}}$ is the $n_{s}\times n_{s}$ identity matrix (see Appendix A.1).

### v2

@@
Abstract
Peer influence on effort devoted to some activity is often studied when effort is unobserved and the researcher instead observes an outcome that combines effort with other shocks. For instance, in education, achievement measures such as GPA reflect both effort and idiosyncratic GPA shocks. We propose an alternative approach that circumvents this approximation. Our framework distinguishes unobserved shocks to GPA that do not affect effort from preference shocks that do affect effort levels. We show that peer effects estimates obtained using our approach can differ significantly from classical estimates (where effort is approximated) if the network includes isolated students. Applying our approach to data on high school students in the United States, we find that peer effect estimates relying on GPA as a proxy for effort are 40% lower than those obtained using our approach.
@@ 1  Introduction
In recent years, there has been a growing interest in understanding how peers impact educational outcomes (sacerdote2011peer; epple2011peer). Because peer effects on students’ academic effort may involve a social multiplier effect, understanding whether students are influenced by their friends and the size of this influence is crucial for evaluating policies aimed at improving academic achievement (m1993). However, estimating peer effects based on academic effort presents challenges, since effort itself is generally unobserved. Consequently, many empirical studies estimate peer effects on achievement (test scores or GPA) as the policy-relevant outcome, often grounding this choice in models where achievement depends on own and peers’ effort. While microfounded models exploring the impact of peer interactions on academic achievement are often focused on effort, GPA is often the main observed achievement measure in the empirical analysis, especially in settings where direct measures of student effort are limited or unavailable (e.g., see cpz2009; arcidiacono2012estimating; fruehwirth2013identifying; fruehwirth2014can; hong2017sitting). Yet, this approximation overlooks that GPA is not solely influenced by effort but also by many other factors, including unobserved student and school characteristics.
@@
In this paper, we investigate the implications of using GPA as a proxy for academic effort on the estimation and interpretation of peer effects. We develop a structural model of educational effort with social interactions, in which students decide on their academic effort while taking into account that of their peers (friends). We also explicitly model the production function of academic achievement (GPA), which includes effort as a key input. Without using a proxy for student academic effort, we show that our model identifies peer effects on academic effort. Furthermore, we demonstrate that the size of the effect may differ from that estimated using GPA as a proxy for effort, because the proxy-based approach suffers from an issue akin to an omitted variables bias. From the structural model, we derive a reduced-form equation for GPA, which differs from the classical linear-in-means peer effects specification, that highlights the need to control for two types of unobserved GPA shocks to disentangle peer effects on academic effort from other common effects captured by GPA. First, one needs to account for common shocks that directly influence GPA, such as improvements in teaching quality, irrespective of the effort level. These shocks result in a GPA increase for the same level of effort and do not involve a social multiplier effect. Second, one needs to account for common shocks affecting students’ preferences, such as increasing motivation to value academic achievement through information, which influence both academic effort and GPA, and may have social multiplier effects among students with friends.
@@
We demonstrate that approximating student effort with GPA may result in biased estimates of peer effects when some students in the network are isolated (i.e., have no friends). This is because students who have friends are affected differently by the two types of GPA shocks mentioned above compared to isolated students. Standard approaches that use GPA as a proxy fail to differentiate between these two types of shocks, leading to biased estimates of peer effects when the network includes isolated students. However, we show that in networks without isolated students, the standard model does not produce biased estimates of peer effects. The key difference between the reduced-form of the standard model and our framework is that while the standard model has a single intercept for each school, our approach accounts for unobserved school-level heterogeneity based on whether a student has friends or not. This amounts to introducing two types of fixed effects per school, one for students with friends and another for isolated students. This two‐fixed‐effect approach is uncommon in linear‐in‐means models—one exception is boucher2024toward, who use it for different identification purposes. [^1] Their analysis requires identifying peer effects from the equation for isolated individuals. As a result, they treat the GPA equation separately for isolated and non-isolated nodes.
@@
Our econometric model incorporates two types of school-level unobserved shocks to GPA, which renders the well-known reflection problem more complex. [^2] The reflection problem arises when one cannot disentangle endogenous peer effects from exogenous contextual peer effects (m1993) In the case of the standard linear-in-means model, BramoulleDjebbariFortin2009 provide straightforward conditions relating to the network structure under which the reflection problem is resolved. We extend their identification analysis to our framework; our main condition for identification requires that the network must include at least two students separated by a path of distance three, which is slightly stronger than the condition in BramoulleDjebbariFortin2009.
We illustrate these points through both a Monte Carlo simulation study and an empirical application using data from the National Longitudinal Study of Adolescent to Adult Health (Add Health). Estimating peer effects through our proposed method suggests that increasing the average GPA of peers by one point leads to a 0.856 point increase in students’ GPA. In contrast, the standard linear-in-means model using GPA as a proxy for student academic effort estimates this effect at 0.507. This substantial difference highlights that failing to account for two types of shocks at the school level may yield substantially biased estimates. Importantly, we also find that other plausible approaches, such as estimating a model that accounts for a single school-level fixed effect and incorporating a dummy variable for isolated students, fail to rectify the bias, as does excluding isolated students from the sample. Finally, we show that the standard model is sensitive to plausible forms of measurement error: when a share of students is misclassified as isolated, its peer effect estimate declines, whereas our double fixed effects model remains stable.
Isolated nodes—students without direct friendship connections—are a common feature in studies examining directed social networks. In the Add Health dataset, 22% of students have no friends, although 11% are not entirely isolated, as they are nominated by others as friends. Several studies using social network data from educational settings, such as alan2021building, conti2013popularity, boucher2021ethnic, banerjee2024changes, report similar patterns. We show that even with a small share of isolated nodes, the bias can be large.
@@
We also extend our analysis to the case of endogenous networks. Network endogeneity can occur because we do not observe certain student characteristics, such as intellectual quotient (IQ), that may influence both students’ likelihood to form links with others and their GPA. We control for these unobserved characteristics using a two-stage estimation approach. Our method is nonparametric as in johnsson2021estimation. We do not impose a specific parametric restriction in the relationship between the unobserved characteristics and GPA. Our approach is similar to generalized additive models (GAM), which are widely employed in the nonparametric regression literature (see hastie2017generalized).
@@
Our structural model and econometric approach can also be applied to study peer influence on other outcomes that depend on exerted effort. An example is the body mass index (BMI), which cannot be directly chosen (e.g., fy2015). People need to exert effort, such as developing healthy diet habits or engaging in physical exercise to improve their BMI. Peer influences are more related to effort than BMI. Another example is peer effects on a worker’s effort (e.g., mas2009peers; cornelissen2017peer). The observed outcome is generally worker’s productivity, whereas peer effects take source in effort.
@@
There is a large literature studying social interactions both theoretically and empirically (durlauf2010social; blume2011identification). We follow the games in networks approach to the analysis of social interactions (see jackson2015games, for a comprehensive overview of this literature). BallesterZenou2006 analyze a noncooperative game with linear-quadratic utilities and strategic complementarities, in which each player decides how much effort they exert. Applying a similar framework to an education application, cpz2009 use GPA as a proxy for the exerted effort. We contribute to this strand of the literature by explicitly modelling the production function that captures how effort, along with other factors, translates into GPA, while distinguishing between unobserved shocks that directly impact GPA without affecting effort, from shocks that impact both academic effort and GPA. This leads to a reduced-form equation for GPA that differs from the standard linear-in-means peer effects specification.
@@
This paper makes a methodological contribution to the extensive empirical literature on peer effects on educational outcomes (sacerdote2011peer; epple2011peer). We show both analytically and through an empirical application using AddHealth data that approximating student effort with GPA may result in biased estimates of peer effects, when some students in the network lack friends. Since isolated students are a common feature in many social network datasets, this finding highlights the limitations of using proxy variables, such as GPA, for estimating peer effects reliably.
@@
Our paper also contributes to the econometric literature on peer effects (de2017econometrics; kline2020econometric). A key challenge in this field is the reflection problem (m1993). A recent wave of papers have addressed this issue by imposing conditions on the network structure (BramoulleDjebbariFortin2009; de2010identification). [^3] Other studies address the reflection problem using group size variation (davezies2009identification; lee2007identification) or imposing restrictions on the error terms (graham2008identifying; rose2017identification). For an overview of this literature, see bramoulle2020peer. Our contribution is to study the reflection problem in a setting where the GPA is influenced by various types of common shocks at the school level, and where students may have no peers. As in BramoulleDjebbariFortin2009, our main identification condition involves the network structure and can be readily tested in empirical applications.
@@
The remainder of the paper is organized as follows. Section 2 presents the microeconomic foundations of the model using a network game in which students decide their academic effort. Section 3 describes the econometric model and addresses the identification and estimation of the parameters. Section 4 presents our empirical analysis using Add Health data. Section LABEL:sec:endo provides an extension of our framework to endogenous networks. Section LABEL:sec:conc concludes this paper.
@@ 2  The Structural Model
We consider $S$ independent schools and denote by $n_{s}$ the number of students in the $s$-th school, $s\in\{1,\dots,S\}$. Let $n$ be the total number of students, that is, $n=\sum_{s=1}^{S}n_{s}$. Each student $i$ in school $s$ has a GPA denoted by $y_{s,i}$ and observable characteristics represented by a $K$-vector $\mathbf{x}_{s,i}$. Students interact in their school through a directed network that can be represented by an $n_{s}\times n_{s}$ adjacency matrix $\mathbf{A}_{s}=[a_{s,ij}]_{ij}$, where $a_{s,ij}$ = 1 if student $j$ is $i^{\prime}s$ friend and $a_{s,ij}=0$ otherwise. We assume that $a_{s,ii}$ = 0 for all $i$ and $s$ so that students cannot interact with themselves. In addition, we only consider within-school interactions; students do not interact with peers in other schools (see cpz2009). We define the social interaction matrix $\mathbf{G}_{s}=[g_{s,ij}]_{ij}$ as the row-normalized adjacent matrix $\mathbf{A}_{s}$, that is, $g_{s,ij}=1/n_{s,i}$ if $j$ is a $i$’s friend and $g_{s,ij}=0$ otherwise, where $n_{s,i}$ is the number of friends of student $i$ within school $s$.
Student $i$ chooses their effort $e_{s,i}$, which in turn affects their GPA. More precisely, GPA is a function of student effort $e_{s,i}$ and a random term $\eta_{s,i}$ that captures unobservable characteristics. [^4] Whether or not the random term $\eta_{s,i}$ is observed by student $i$ or their peers is inconsequential for the analysis of the game. Following fruehwirth2013identifying and BoucherFortin2016, we posit that this relationship is given by the production function:
$y_{s,i}=\alpha_{s}+e_{s,i}+\eta_{s,i},$ (1)
where $\alpha_{s}$ captures unobserved school-level GPA shifters, such as teacher quality, operating as fixed effects. The linear production function is somewhat restrictive and can be generalized to a nonlinear or nonparametric production function (e.g., see fruehwirth2013identifying). However, identification of the resulting econometric model can be intractable, especially when the model allows for unobserved school heterogeneity. We also consider the production function $y_{s,i}=\alpha_{s}+\delta e_{s,i}+\mathbf{x}^{\prime}_{s,i}\boldsymbol{\theta}+\eta_{s,i}$, where $\delta>0$ and $\boldsymbol{\theta}$ is a vector of parameters. In Appendix LABEL:append:GPA, we show that both specifications imply the same reduced-form structure for the achievement equation we estimate below, although in the second case, the coefficients on own and contextual covariates should be interpreted as reduced-form composites. We discuss identification in Section 3.1. For parsimony, we retain (1) in the main text and present the alternative specification in Appendix LABEL:append:GPA.
The effort exerted and the GPA obtained provide students with a benefit that is captured by a payoff function, which, as in cpz2009 and BoucherFortin2016, takes a linear-quadratic form:
@@
where $\mathbf{X}_{s}=(\mathbf{x}_{s,1},\dots,\mathbf{x}_{s,n_{s}})^{\prime}$, $\mathbf{g}_{s,i}$ is the $i$-th row of $\mathbf{G}_{s}$, $\mathbf{e}_{s,-i}=(e_{s,1},\dots,e_{s,i-1},e_{s,i+1},\dots,e_{s,n})^{\prime}$, $\mathbf{e}_{s}=(e_{s,1},\dots,e_{s,n_{s}})^{\prime}$, the term $\mathbf{g}_{s,i}\mathbf{e}_{s}$ is the average effort of peers, and $c_{s}$, $\boldsymbol{\beta}$, $\boldsymbol{\gamma}$ are unknown parameters. The parameter $\lambda$ captures endogenous peer effects. The payoff function (2) encompasses two components: a private sub-payoff and a social sub-payoff. The term $(c_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\varepsilon_{s,i})$ represents the benefit enjoyed per unit of GPA achieved, where $\varepsilon_{s,i}$ is the student type (observable by all students). This benefit accounts for student observed heterogeneity, as it depends on $\mathbf{x}_{s,i}$ and peer group average characteristics $\mathbf{g}_{s,i}\mathbf{X}_{s}$ termed contextual variables (see m1993). The benefit also accounts for school unobserved heterogeneity through the parameter $c_{s}$. The second term of the private sub-payoff, $e_{s,i}^{2}/2$ reflects the cost of exerting effort. The social sub-payoff $\lambda e_{s,i}\mathbf{g}_{s,i}\mathbf{e}_{s}$ implies that an increase in the average peer group’s effort $\mathbf{g}_{s,i}\mathbf{e}_{s}$ influences student $i$’s marginal payoff if $\lambda\neq 0$. When $\lambda>0$, the payoff function (2) implies complementarity between students’ and peers’ efforts, whereas $\lambda<0$ indicates substitutability in efforts. [^5] An alternative specification considers the social-payoff as $\frac{\lambda}{2}(e_{s,i}-\mathbf{g}_{s,i}\mathbf{e}_{s})^{2}$, which represents a social cost depending on the gap between the student’s effort level and the average peer effort. This specification leads to conformist preferences when $\lambda>0$. Our approach can also be extended to this alternative specification.
@@
The parameters $\alpha_{s}$ and $c_{s}$ capture different unobserved shocks at the school level, and are conceptually different in terms of their policy implications. $\alpha_{s}$ captures unobserved shocks on GPA that do not affect student effort. These shocks, such as variation in teaching quality and school management, directly impact GPA irrespective of student effort. [^6] This reasoning holds because GPA is unbounded. An increase in $\alpha_{s}$ necessarily results in a higher GPA, and thus a higher payoff, regardless of the effort level. If we were to consider a framework where GPA is bounded, an increase in $\alpha_{s}$ may decrease the effort for students nearing the upper bound of GPA (e.g., see fruehwirth2013identifying). We address this concern directly by re-estimating our main specifications using a Tobit model in Appendix LABEL:append:app (Table LABEL:tab:estexotobit). The ordering of peer-effects estimates across models is unchanged. On the other hand, $c_{s}$ captures shocks on student preferences, particularly on the marginal payoff. For instance, interventions aimed at making students more aware of the returns to academic achievement could influence the marginal payoff. Such a shock can be captured by $c_{s}$ and would influence effort and consequently GPA through equation (1). We show that $\alpha_{s}$ and $c_{s}$ do not impact GPA in the same way (see Section 3.1).
@@
By substituting $y_{s,i}$ from Equation (1) into Equation (2), we obtain a payoff function that does not depend on GPA (see Appendix LABEL:append:NE). This new payoff function defines a static game with complete information, in which students simultaneously choose their effort levels to maximize their payoff. The best response function for the students is given by:
$e_{s,i}=c_{s}+\lambda\mathbf{g}_{s,i}\mathbf{e}_{s}+\mathbf{x}_{s,i}^{\prime}\boldsymbol{\beta}+\mathbf{g}_{s,i}\mathbf{X}_{s}\boldsymbol{\gamma}+\varepsilon_{s,i}.$ (3)
@@
The best response function (3) in matrix form can be expressed as $\mathbf{e}_{s}=c_{s}\mathbf{1}_{n_{s}}+\lambda\mathbf{G}_{s}\mathbf{e}_{s}+\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\boldsymbol{\varepsilon}_{s}$, where $\boldsymbol{\varepsilon}_{s}=(\varepsilon_{s,1},\dots,\varepsilon_{s,n_{s}})^{\prime}$ and $\mathbf{1}_{n_{s}}$ is an $n_{s}$-vector of ones. A solution of this equation in $\mathbf{e}_{s}$ is a Nash equilibrium (NE) of the game. As $\mathbf{G}_{s}$ is row-normalized, the NE is unique under Assumption 2.1 and can be expressed as $\mathbf{e}_{s}=(\mathbf{I}_{n_{s}}-\lambda\mathbf{G}_{s})^{-1}(c_{s}\mathbf{1}_{n_{s}}+\mathbf{X}_{s}\boldsymbol{\beta}+\mathbf{G}_{s}\mathbf{X}_{s}\boldsymbol{\gamma}+\boldsymbol{\varepsilon}_{s})$, where $\mathbf{I}_{n_{s}}$ is the $n_{s}\times n_{s}$ identity matrix (see Appendix LABEL:append:NE).